OpenStudy (anonymous):

can someone help please i want to be sure i did this right. question? During one year, the Peddy's real estate bill included $948 for county improvements. Of this amount, $308 went for transportation and roads. What percent did the Peddy family pay for transportation and roads?

OpenStudy (lgbasallote):

\[308 \div 948 \times 100\] is this what you did?

OpenStudy (anonymous):

no all i did was divide the two

OpenStudy (lgbasallote):

why 2?

OpenStudy (anonymous):

sorry i mean 308 / 948 I divided these two numbers

OpenStudy (lgbasallote):

ohhh..yeah that's right...but you need to multiply by 100 to get the percent

OpenStudy (lgbasallote):

if you dont multiply by 100 you just get the decimal

OpenStudy (lgbasallote):

got it?

OpenStudy (anonymous):

yes thank you

OpenStudy (lgbasallote):

welcome :D

OpenStudy (anonymous):

can help with another one

OpenStudy (anonymous):

The second angle of a triangle is 3 times as large as the first. The third angle is 55 degrees more then the first. Find all three angles.

OpenStudy (anonymous):

I got 25 degrees for the first angle, 75 degrees for the second angle and 80 degree for the third angle. is this right

OpenStudy (lgbasallote):

is this a triangle?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

let x = first angle y = second angle z = third angle y = 3x z = x + 55 x + y + z = 180 x + 3x +x + 55 = 180 5x + 55 = 180 5x = 125 x = 25 y = 25 x 3 = 75 z = 25 + 55 = 80 you're right

OpenStudy (anonymous):

ty i have more if you are available to help

OpenStudy (lgbasallote):

uhhh...you really should just post it as new questions so more can help =)) i'll jump in if i know how to do your questions

OpenStudy (anonymous):

ok ty